# Classification of Cellular Automata based on the Hamming distance

### v1

@@ Abstract
Elementary cellular automata are the simplest form of cellular automata, studied extensively by Wolfram in the 1980s. He discovered complex behavior in some of these automata and developed a classification for all cellular automata based on their phenomenology. In this paper, we present an algorithm to classify them more effectively by measuring difference patterns using the Hamming distance. Our classification aligns with Wolfram’s and further categorizes them into additional subclasses.
@@
preprint: APS/123-QED
@@ I Introduction
Cellular automata (CA) [1, 2] are systems that can be used to produce simple agent-based models exhibiting complex dynamics. Each cell can exist in several possible states and is updated at each time iteration based on the state of neighboring cells, its own state, and one or several update rules. They are employed in various contexts and have broad applications in physics [3, 4, 5, 6, 7], engineering [8], cryptography [9, 10], and theoretical biology [11, 12, 13].
@@ II Evolutionary Cellular Automata and Wolfram’s Classification
We present the Wolfram classification, compiled from WolframAlpha [^1] WolframAlpha computational engine of knowledge https://www.wolframalpha.com in Table 2 in the Appendix alongside our own classification.

### v2

@@ Abstract
Elementary cellular automata are the simplest form of cellular automata, studied extensively by Wolfram in the 1980s. He discovered complex behavior in some of these automata and developed a classification for all cellular automata based on their phenomenology. In this paper, we present an algorithm to classify them more effectively by measuring difference patterns using the Hamming distance. Our classification aligns with Wolfram’s and further categorizes them into additional subclasses. Finally, we have found a heuristic reasoning providing and explanation about why some rules evolve into fractal patterns.
@@
preprint: AIP/123-QED
The classification of cellular automata has been a prevalent topic in the field. Wolfram’s classification significantly advanced the identification of the main dynamical differences between various automata. Four different types of dynamical behavior identify the classes of automata whether the system is at equilibrium, is periodic, chaotic or present dissipative and transient complex behavior. Traditionally, determining the class of a rule requires analyzing the space-time dynamics and making a subjective judgment. The difference patterns between two configurations, where an initial cell differs in one configuration, also exhibit distinct dynamical behavior for each class, but the assessment remains subjective. We have discovered that plotting the Hamming distance between the two configurations (i.e., the number of differing cells) over time provides an objective and quicker method for classifying automata into the same four Wolfram classes. If the distance is null, constant or periodic, complex, or shows transient chaos, then the rule belongs to Classes 1 through 4, respectively. Furthermore, we have found a heuristic reasoning explaining the fractal nature of some rules.
@@ I Introduction
Cellular automata (CA) [1, 2] are systems that can be used to produce simple agent-based models exhibiting complex dynamics. Each cell can exist in several possible states and is updated at each time iteration based on the state of neighboring cells, its own state, and one or several update rules. They are employed in various contexts and have broad applications in physics [3, 4, 5, 6, 7, 8, 9], engineering [10], cryptography [11, 12], and theoretical biology [13, 14, 15].
@@
Wolfram’s classification is extensively used today as in [17], for example, and other classifications have been developed so far [18, 19, 20, 21, 22, 23, 24]. The authors of [25] claim that most classification are limited to ECAs. Wolfram’s classification, as well as our own, were studied under ECAs but are extensive to other CAs.
@@ II Evolutionary Cellular Automata and Wolfram’s Classification
We present the Wolfram classification, compiled from WolframAlpha [26] in Table 2 in the Appendix alongside our own classification.